OpenStudy (konradzuse):

Use the tabular method to find the integral. 5x^3 sin(x)dx

OpenStudy (konradzuse):

u = x^3 or 5x^3 du = 3x^2 or 15x^2 I believe I can pull the 5 out tho? dv = sin(x)dx v = -cos(x)

OpenStudy (konradzuse):

\[5x^3* -\cos(x) - \int\limits -\cos(x)15x^2dx\]

OpenStudy (konradzuse):

u = 15x^2 du = 30xdx dv = -cos(x)dx v = -sin(x)

OpenStudy (konradzuse):

\[5x^3* -\cos(x) + \sin(x)30xdx + \int\limits \sin(x) 30xdx\]

OpenStudy (konradzuse):

love how openstudy just crashed had rto re-write all of that :P

OpenStudy (konradzuse):

now 1 more time u = 30x du = 30dx dv = sin(x)dx v = -cos(x)

OpenStudy (anonymous):

Continuously tabular until the third time: = 5 ( - x³cosx + 3x² sinx + 6x cosx - 6sinx ) + C

OpenStudy (konradzuse):

\[5x^3 *-\cos(x) + \sin(x) 30x + 30x * -\cos(x) + 30 \int\limits \cos(x)dx\]

OpenStudy (konradzuse):

???

OpenStudy (konradzuse):

\[5x^3 *-\cos(x) + 15x^2\sin(x) + 30x * -\cos(x) + 30 \sin(x) +c\]

OpenStudy (anonymous):

You should learn Tabular method!

OpenStudy (konradzuse):

did I not do it correctly?

OpenStudy (konradzuse):

seems to be the same answer as yours besides some signage issues.

OpenStudy (anonymous):

Your final one is correct, but your method isn't Tabulation which plainly simplify!

OpenStudy (anonymous):

That's why you got wrong signs!

OpenStudy (konradzuse):

yeah I see it now, creating a table.... I realized when I was doing it that if I said it to be +sin(x) it wou;dn't have cancelled, but I left it as -sin(x) so it did... weird.

OpenStudy (konradzuse):

Ic.. hmm oh wells :P.

OpenStudy (konradzuse):

\[-5x^3\cos(x) + 15x^2\sin(x) +30xcos(x) - 30\sin(x) +c\]

OpenStudy (anonymous):

I just multiply in, see if it matches with yours: = -5 x³cosx + 15x² sinx + 30 x cosx - 30 sinx ) + C

OpenStudy (konradzuse):

mhm